OpenStudy (kearsten):

solve the inequality, 2x+3y ≤6 show steps plz

OpenStudy (michele_laino):

here we can only solve your inequality, graphically

OpenStudy (michele_laino):

in order to that, first we have to draw the line whose equation is: 2x+3y=6

OpenStudy (michele_laino):

I will show all needed steps

OpenStudy (michele_laino):

Let's consider the equation: 2x+3y=6

OpenStudy (michele_laino):

If I set x=0, I can write: 2*0+3y=6 then what is y?

OpenStudy (kearsten):

2?

OpenStudy (michele_laino):

that's right! So we know that our line, will pass at point (0,2)

OpenStudy (michele_laino):

Now, if I set y=0, I can write: 2x+3*0=6 what is x?

OpenStudy (kearsten):

3?

OpenStudy (michele_laino):

ok! so our line also passes at point (3,0)

OpenStudy (michele_laino):

here is your line: 

OpenStudy (kearsten):

thanks for the help

OpenStudy (michele_laino):

now if I solve your inequality for y, I get: \[\Large y \leqslant 2 - \frac{2}{3}x\]

OpenStudy (michele_laino):